If"\ f(x) = 3x^2 – 2x + 5" , find f[1, 2], f[2, 3] and f[1, 2, 3].


1
Expert's answer
2021-07-20T17:18:50-0400

The function "f(x)=3x^2-2x+5" is continuous for "x\in \R" as polynomial.



"f'(x)=6x-2"

"f'(x)=0=>6x-2=0=>x=\dfrac{1}{3}"

If "x>\dfrac{1}{3}, f'(x)>0, f(x)" is strictly increasing.



"f(1)=3(1)^2-2(1)+5=6"

"f(2)=3(2)^2-2(2)+5=13"

"f(3)=3(3)^2-2(3)+5=26"

"f:[1,2]\to[6,26]"

"f:[2,3]\to[13,26]"

"f:\{1,2,3\}\to\{6, 13,26\}"